# SADA: Safe and Adaptive Inference with Multiple Black-Box Predictions

### v1

@@ 4.1 Properties of the SADA estimator
(i) The parameter $\bm{\theta}$$^{*}$ lies in the interior of $\Theta$, and $\Theta$ is compact in $\mathbb{R}^{p}$; (ii) $\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})\}\neq 0$ if ${\bm{\theta}}\neq{\bm{\theta}}^{*}$; (iii) $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})$ is differentiable with respect to $\bm{\theta}$} in a neighborhood of $\bm{\theta}$$^{*}$, and $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$ exists and nonsingular. (iv) $\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})$ and $\mathbf{s}(\mathbf{X},\widehat{Y}_{k};{\bm{\theta}})$ have bounded first and second order moments in a neighborhood of $\bm{\theta}$$^{*}$ for $k=1,\dots,K$; (v) $\widehat{\mathcal{W}}^{\text{opt}}\xrightarrow{p}\mathcal{W}^{\text{opt}}$.
@@
The proof of Theorem 1 is provided in Appendix E. In the general case, Theorem 1 shows that the SADA estimator enjoys a guaranteed efficiency gain over the naive estimator, regardless of the quality of $\widehat{\mathbf{Y}}$. This ensures valid and more informative inference for $\bm{\theta}$$^{*}$; see Appendix B for details. The following theorem formalizes the adaptivity property of the SADA estimator and is proved in Appendix F.
@@
5 Experiments
5.1 Synthetic datasets
@@
We first conduct small-scale simulations to evaluate the performance of the proposed algorithm. We generate $Y\sim N(\theta^{*},1)$, and generate two predictions $\widehat{Y}_{1}=\gamma Y+(1-\gamma)\epsilon_{1}$ and $\widehat{Y}_{2}=(1-\gamma)Y+\gamma\epsilon_{2}$, where $\gamma$ is a tuning parameter and $\epsilon_{1},\epsilon_{2}$ are independent white noises drawn from the standard normal distribution. The true parameter $\theta^{*}=0.5$. As $\gamma$ increases from 0 to 1, the quality of the prediction $\widehat{Y}_{1}$ improves in approximating the ground truth $Y$, while the prediction $\widehat{Y}_{2}$ becomes less accurate. We generate a dataset of size $N=200$, among which $n=60$ are labeled. We perform 1000 Monte Carlo replications.
Figure 2 reports the relative efficiency of different methods compared to the naive method, calculated as the ratio of their standard deviations across replications, as $\gamma$ increases from 0 to 1. Figure 2(a) shows that the performance of the PPI estimator heavily depends on the accuracy of the prediction—it performs even worse than the naive estimator when the prediction is poor. Figure 2(b) demonstrates that PPI++ provides protection against poor prediction quality, yielding a variance that is never greater than that of the naive method. It reduces to the naive method when the prediction is noninformative. Figure 2(c) illustrates that our proposed SADA estimator not only consistently maintains a standard deviation lower than that of the naive estimator but also adaptively combines the strengths of $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ in their respective regions, regardless of which performs better, resulting in the most stable overall performance.
@@
Figure 2: Relative efficiency of different methods compared to the naive method under varying prediction quality. Stars and triangles on the line indicate standard deviations over 1000 replications; scatter points represent those from individual replications.
5.2 Benchmark datasets
Wine reviews
In this section, we apply our proposed methods to the wine review data published on WineEnthusiast: https://www.kaggle.com/datasets/mysarahmadbhat/wine-tasting. The dataset includes characteristics of the wine such as country, price, and region, along with reviews and ratings (from 80 to 100) from wine tasters. The goal is to estimate the mean rating. We use three predictive models–GPT-4o, Llama-3-8B, and DeepSeek–to generate predictions of the ratings based on the characteristics and the taster’s reviews. Appendix G provides a step-by-step guidance of how to generate predictions using these LLMs. We sample 5000 instances for our analysis, using 1500 of them as unlabeled data and varying the number of labeled samples $n$ from 600 to 3500. The ratings of all 5000 instances are treated as the ground truth.
Figure 3 reports the variation of standard deviations of different methods as the labeled data size increases. The PPI and PPI++ methods in the subfigures are implemented using different single LLMs. While the standard deviations of all methods generally decrease with more labeled data, the PPI estimator may perform worse than the naive method. In contrast, PPI++ guarantees improved efficiency over the naive approach, but the gains are limited when prediction accuracy is low, as seen in Figure 3(b). The SADA estimator, which integrates all predictions, consistently outperforms both PPI and PPI++ methods. Notably, its performance closely matches the performance of the best-performing PPI++ estimator using GPT-4o, highlighting its adaptive nature.
We also plot the oracle SADA estimator, SADAo, which includes the ground truth as one of its predictions. As shown, SADAo performs almost identically to the oracle estimator, calculated as the mean of the ground truth. This result verifies the desired adaptivity of the proposed method.
Figure 3: Comparison of standard deviations of different methods leveraging various prediction strategies. The estimand is the mean rating in the wine reviews dataset.
Politeness of online requests
We also apply our method to study the relationship between politeness and the use of indicative modal features, using the dataset from Danescu-Niculescu-MizilSudhofJurafskyLeskovecEtAl2013. From this dataset, we select 1000 requests from Stack Exchange and Wikipedia, each rated on a politeness scale from 1 to 25, averaged across five human annotators. The parameter of interest is the regression coefficient obtained by regressing the politeness score on the indicative modal features. We randomly designate 300 of the requests as unlabeled data and increase the number of labeled samples $n$ from 50 to 700.
Figure 4 shows similar results as Figure 3. In this task, the PPI estimator performs significantly worse than the naive method when predictions are inaccurate, as seen in Figure 4(a)-(b). Notably, PPI++ does not guarantee a lower variance than the naive approach in this setting. In contrast, SADA improves upon all other methods under different sample sizes.

### v2

@@ 3.2 Theoretical guarantees
(i) The parameter $\bm{\theta}$$^{*}$ lies in the interior of $\Theta$, and $\Theta$ is compact in $\mathbb{R}^{p}$; (ii) $\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})\}\neq 0$ if ${\bm{\theta}}\neq{\bm{\theta}}^{*}$; (iii) $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})$ is differentiable with respect to $\bm{\theta}$} in a neighborhood of $\bm{\theta}$$^{*}$, and $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$ exists and nonsingular; (iv) $\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})$ and $\mathbf{s}(\mathbf{X},\widehat{Y}_{k};{\bm{\theta}})$ have bounded first and second order moments in a neighborhood of $\bm{\theta}$$^{*}$ for $k=1,\dots,K$; (v) $\widehat{\mathcal{W}}^{\text{opt}}\xrightarrow{p}\mathcal{W}^{\text{opt}}$.
The following theorem formalizes the safety property of the SADA estimator and is proved in Supplementary Section S5.
@@
Theorem 1 shows that the SADA estimator enjoys a guaranteed efficiency gain over the naive estimator, regardless of the quality of $\widehat{\mathbf{Y}}$. This ensures valid and more informative inference for $\bm{\theta}$$^{*}$; see Supplementary Section S2 for details. The following theorem formalizes the adaptivity property and is proved in Supplementary Section S6.
@@
Theorem 2(i) shows that, in the oracle scenario where one of the predictions coincides with the ground truth outcome, the SADA estimator effectively increases the sample size from $n$ to $N$ and thereby attains a faster convergence rate. Theorem 2(ii) shows that, when all predictions are functions of the covariates $\mathbf{X}$ and one score function equals the corresponding conditional mean, the estimator attains the semiparametric efficiency bound. Importantly, we never need to know in advance which prediction is the best; the entire procedure is data driven, which makes it appealing in practice.
4 SADA for prediction
The previous section focuses on estimation and inference for the parameter $\bm{\theta}$$^{*}$, where the primary goal is to obtain an unbiased and more efficient estimator. In other applications, such as image classification, practitioners are primarily concerned with predictive accuracy rather than parameter estimation itself, and in such cases the SADA framework needs to be adapted accordingly. Intuitively, even if one estimator, say, $\widehat{\bm{\theta}}_{1}$, is closer to $\bm{\theta}$$^{*}$ than the other estimator $\widehat{\bm{\theta}}_{2}$ in mean squared error, it does not in general imply a smaller predictive risk; that is, $\mathcal{R}(\widehat{\bm{\theta}}_{1})\leq\mathcal{R}(\widehat{\bm{\theta}}_{2})$. Motivated by this distinction, in this section we slightly modify the SADA procedure to explicitly target minimization of the predictive risk $\mathcal{R}({\bm{\theta}})$ and discuss how it can be designed to achieve better prediction performance.
To build intuition, consider an unbiased estimator $\widehat{\mathcal{R}}({\bm{\theta}})$ of $\mathcal{R}({\bm{\theta}})$ such that $\mathbb{E}\{\widehat{\mathcal{R}}({\bm{\theta}})\}=\mathcal{R}({\bm{\theta}})$. Define $\widehat{\bm{\theta}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}({\bm{\theta}})$. Then, in a heuristic sense, the excess risk $\mathcal{R}(\widehat{\bm{\theta}})-\mathcal{R}({\bm{\theta}}^{*})\leq\mathcal{R}(\widehat{\bm{\theta}})-\widehat{\mathcal{R}}(\widehat{\bm{\theta}})+\widehat{\mathcal{R}}({\bm{\theta}}^{*})-\mathcal{R}({\bm{\theta}}^{*})\leq 2\sup_{{\bm{\theta}}\in\Theta}|\widehat{\mathcal{R}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})|$, where the first inequality follows from $\widehat{\mathcal{R}}(\widehat{\bm{\theta}})\leq\widehat{\mathcal{R}}({\bm{\theta}}^{*})$ by definition. Under some regularity conditions, standard uniform concentration results indicate that this supremum depends on the complexity of $\Theta$, the variance scale of $\widehat{\mathcal{R}}$, and the tail behavior of the empirical loss. Accordingly, once class complexity is fixed, reducing the variance of $\widehat{\mathcal{R}}$ is a direct and effective way to tighten the uniform deviation and lower the excess risk.
Let $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})=n^{-1}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})$ denote the naive empirical loss function which only uses the labeled data, and $\widehat{\bm{\theta}}^{\text{nv}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$. We consider a set of empirical loss functions, indexed by $\bm{\omega}=(w_{1},w_{2},\dots,w_{K})^{\mathrm{\scriptscriptstyle T}}$,
$\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})=\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})+\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})\right\},$ (7)
where $\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),\widehat{\mathbf{y}})=(\ell(f_{\bm{\theta}}(\mathbf{x}),\widehat{y}_{1}),\ell(f_{\bm{\theta}}(\mathbf{x}),\widehat{y}_{2}),\dots,\ell(f_{\bm{\theta}}(\mathbf{x}),\widehat{y}_{K}))^{\mathrm{\scriptscriptstyle T}}$. Noting that the second term has zero expectation for any $\bm{\omega}$ and any distribution of $\widehat{\mathbf{y}}$, therefore, equation (7) always defines an unbiased estimator for $\mathcal{R}({\bm{\theta}})$. As a special case, when $\bm{\omega}=\mathbf{0}$, $\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})$ reduces to $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$. For a fixed $\bm{\theta}$}, we propose to select the optimal loss function from the set by minimizing the mean squared error between $\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})$ and the expected loss $\mathcal{R}({\bm{\theta}})$. We have
$\bm{\omega}^{\text{opt}}({\bm{\theta}})=$ $~\operatornamewithlimits{arg\,min}_{\bm{\omega}}\mathbb{E}\{\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})-\mathcal{R}({\bm{\theta}})\}^{2}$
$=$ $~\frac{N-n}{N}\operatorname{var}\{\mathcal{L}(f_{\bm{\theta}}(\mathbf{X}),\widehat{\mathbf{Y}})\}^{-1}\operatorname{cov}\{\mathcal{L}(f_{\bm{\theta}}(\mathbf{X}),\widehat{\mathbf{Y}}),\ell(f_{\bm{\theta}}(\mathbf{X}),Y)\}.$
Then, we define $\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\widehat{\mathcal{R}}({\bm{\theta}},\widehat{\bm{\omega}}^{\text{opt}}({\bm{\theta}}))$, where $\widehat{\bm{\omega}}^{\text{opt}}$ is a consistent estimator of $\bm{\omega}^{\text{opt}}$, and obtain the SADA estimator as $\widehat{\bm{\theta}}^{\text{sada}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})$. It is worth noting that the optimal weight depends on $\bm{\theta}$} and, in principle, should be optimized jointly with the parameter within the loss function, which may lead to numerical difficulties and instability in practice. Therefore, for implementation, we recommend an iterative procedure that alternately updates the weight $\bm{\omega}^{\text{opt}}$ and the target parameter $\bm{\theta}$$^{*}$, as outlined in Algorithm 1.
Algorithm 1 SADA for prediction
1:Input: Observations $(\mathbf{x}_{i},y_{i},\widehat{\mathbf{y}}_{i})_{i=1}^{n}\cup(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i})_{i=n+1}^{N}$, prediction function $f_{\bm{\theta}}$, loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$, number of iterations $J$.
2: Compute the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}}n^{-1}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})$.
3: Initialize $\widehat{\bm{\theta}}^{(0)}=\widehat{\bm{\theta}}^{\text{nv}}$.
4:for $j=1,\ldots,J$ do
5:  Compute the optimal weight
$\widehat{\bm{\omega}}^{(j)}=\frac{N-n}{N}$ $\Bigg\{\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}\right)\mathcal{L}\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}\right)^{\mathrm{\scriptscriptstyle T}}\Bigg\}^{-1}$
$~~\times\Bigg\{\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}\right)\ell\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),y_{i}\right)\Bigg\}.$
6:  Update the parameter estimate
$\widehat{\bm{\theta}}^{(j)}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})+\Bigg\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})\Bigg\}^{\mathrm{\scriptscriptstyle T}}\widehat{\bm{\omega}}^{(j)}.$
7:end for
8:Output: SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}=\widehat{\bm{\theta}}^{(J)}$.
4.1 Theoretical properties
In this section, we investigate the theoretical properties of $\widehat{\bm{\theta}}^{\text{sada}}$ in terms of its excess risk, $\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$. We first introduce some notations. For a metric space $(\Theta,d)$, let $N(\epsilon,\Theta,d)$ denote the covering number, i.e., the minimum number of balls of radius $\epsilon$ with respect to the metric $d$ that cover the space $\Theta$. For a positive constant $C$, define a class of sub-Gaussian random variables
$\mathcal{F}_{C}:=\left\{\xi:\text{$\xi$ is $\sigma$-sub-Gaussian for some }\sigma\leq C\sqrt{\operatorname{var}(\xi)}\right\}.$
We use some simplified notations when no confusion arises. Let $\ell_{\bm{\theta}}=\ell(f_{\bm{\theta}}(\mathbf{X}),\mathbf{Y})$, $\widetilde{\ell}_{\bm{\theta}}=\ell_{\bm{\theta}}-\mathcal{L}(f_{\bm{\theta}}(\mathbf{X}),\widehat{\mathbf{Y}})^{\mathrm{\scriptscriptstyle T}}\bm{\omega}^{\text{opt}}({\bm{\theta}})$, and $\widehat{\ell}_{\bm{\theta}}=N/(N-n)(\ell_{\bm{\theta}}-\widetilde{\ell}_{\bm{\theta}})$. We endow the parameter space $\Theta$ with three metrics: $d_{1}({\bm{\theta}},{\bm{\theta}}^{\prime})=\operatorname{sd}(\ell_{\bm{\theta}}-\ell_{{\bm{\theta}}^{\prime}})$, $d_{2}({\bm{\theta}},{\bm{\theta}}^{\prime})=\operatorname{sd}(\widehat{\ell}_{\bm{\theta}}-\widehat{\ell}_{{\bm{\theta}}^{\prime}})$ and $d_{3}({\bm{\theta}},{\bm{\theta}}^{\prime})=\operatorname{sd}(\widetilde{\ell}_{\bm{\theta}}-\widetilde{\ell}_{{\bm{\theta}}^{\prime}})$, where $\operatorname{sd}(U)=\sqrt{\operatorname{var}(U)}$ denotes the standard deviation for a random variable $U$. We impose the following regularity conditions, which control the smoothness and tail behavior of the loss functions and the complexity of the parameter space.
Assumption 2.
(i) The loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$ is continuous in $\bm{\theta}$} and uniformly bounded for all ${\bm{\theta}}\in\Theta$ and $(x,y)\in\mathcal{X}\times\mathcal{Y}$; (ii) There exists a positive constant $C>0$ such that, $\forall{\bm{\theta}},{\bm{\theta}}^{\prime}\in\Theta$, $\ell_{\bm{\theta}}-\ell_{{\bm{\theta}}^{\prime}}$, $\widehat{\ell}_{\bm{\theta}}-\widehat{\ell}_{{\bm{\theta}}^{\prime}}$ and $\widetilde{\ell}_{\bm{\theta}}-\widetilde{\ell}_{{\bm{\theta}}^{\prime}}$ all belong to $\mathcal{F}_{C}$. (iii) The Dudley’s entropy integral is finite for $d_{1},d_{2}$ and $d_{3}$; that is, for $j=1,2,3,$
$\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{j})}d\epsilon<\infty.$
Theorem 3 (Excess risk).
Under Assumption 2, we have that with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})\leq\left\{\begin{aligned} &\mathcal{O}\left(\sqrt{\frac{1}{N}\log\frac{1}{\delta}}+\sqrt{\frac{1}{N}}\right),~~\text{if $\widehat{Y}_{k}=Y$ for some $k\in\{1,\dots,K\}$},\\ &\mathcal{O}\left(\sqrt{\frac{1}{n}\log\frac{1}{\delta}}+\sqrt{\frac{1}{n}}+\sqrt{\frac{1}{N}\log\frac{1}{\delta}}+\sqrt{\frac{1}{N}}\right),~~\text{otherwise.}\end{aligned}\right$.
A more precise non-asymptotic bound for $\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$ is given in Theorem S4 in Supplementary Section S7. As a comparison, Theorem S4 also shows that the excess risk of the naive estimator, $\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})$, converges at the order $\mathcal{O}(\sqrt{n^{-1}\log(\delta^{-1})}+\sqrt{n^{-1}})$ with high probability. This implies that the excess risk of the SADA estimator converges at a rate no slower than that of the labeled data only baseline, regardless of the quality of the predictions. In contrast, if one of the predictions perfectly fits the ground truth, although it is unknown which one, the SADA estimator achieves an even faster convergence rate, effectively increasing the sample size from $n$ to $N$.
5 Simulation studies
@@
In this section, we conduct some simulation studies to evaluate the performance of the proposed method. We generate $Y\sim\mathcal{N}(\theta^{*},1)$ with $\theta^{*}=0.5$, and generate two predictions $\widehat{Y}_{1}=\gamma Y+(1-\gamma)\epsilon_{1}$ and $\widehat{Y}_{2}=(1-\gamma)Y+\gamma\epsilon_{2}$, where $\gamma$ is a tuning parameter and $\epsilon_{1},\epsilon_{2}$ are independent white noises drawn from the standard normal distribution. The goal is to estimate and to conduct inference for $\theta^{*}$. As $\gamma$ increases from 0 to 1, the quality of the prediction $\widehat{Y}_{1}$ improves in approximating the ground truth $Y$, while the prediction $\widehat{Y}_{2}$ becomes less accurate. We generate a dataset of size $N=200$, among which $n=60$ observations are labeled, and repeat the experiment over 1000 Monte Carlo replications.
We include naive, PPI, and PPI++ methods for comparison. Table 2 reports the relative efficiency of each method relative to the naive method, defined as the ratio of their standard deviations across replications as $\gamma$ varies from 0 to 1, and Figure 2 presents the corresponding plots. The results on bias and coverage probability are reported in Supplementary Section S8, as they show negligible differences across methods and are not the primary criteria for assessing their performance. It is shown that, the performance of the PPI estimator is highly sensitive to prediction accuracy: it can perform even worse than the naive estimator when the prediction is poor, as illustrated in Figure 2(a), where the ratios of standard deviations exceed 1 when $\gamma$ approaches 0 for PPI using $\widehat{Y}_{1}$ and 1 for PPI using $\widehat{Y}_{2}$. By contrast, PPI++ safeguards against low quality predictions, achieving a variance that is never larger than that of the naive estimator, as shown in Figure 2(b). However, its efficiency gains arise only when the predictions are informative, and it reduces to the naive estimator when the predictions are noninformative.
In comparison, the SADA estimator exhibits several favorable properties. First, it consistently attains a smaller standard deviation than the naive estimator across the entire range of prediction quality. Second, it adaptively exploits the relative strengths of $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ in the regions where each is more informative, which leads to the most stable performance among all methods considered. Third, it is able to incorporate information from both predictions through data-driven weights. This is particularly evident at $\gamma=0.5$, where SADA has substantially lower standard deviations than PPI++ under either prediction.
Table 2: Relative efficiency (ratio of standard deviations for each method to that of the naive method) under varying prediction quality $\gamma$.
Methods Relative efficiency
$\gamma=0$ $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.4$ $\gamma=0.5$ $\gamma=0.6$ $\gamma=0.7$ $\gamma=0.8$ $\gamma=0.9$ $\gamma=1$
PPI $\widehat{Y}_{1}$ 1.549 1.396 1.247 1.104 0.970 0.842 0.736 0.653 0.608 0.610 0.654
$\widehat{Y}_{2}$ 0.655 0.609 0.611 0.656 0.737 0.845 0.968 1.103 1.245 1.392 1.551
PPI++ $\widehat{Y}_{1}$ 0.989 0.982 0.965 0.933 0.877 0.799 0.713 0.636 0.582 0.556 0.549
$\widehat{Y}_{2}$ 0.549 0.556 0.585 0.639 0.715 0.801 0.876 0.932 0.965 0.982 0.989
SADA both 0.548 0.555 0.584 0.636 0.697 0.727 0.697 0.635 0.583 0.555 0.548
@@
Figure 2: Relative efficiency of different methods compared to the naive method as prediction quality varies. Stars and triangles on the line indicate standard deviations over 1000 replications; scatter points represent those from individual replications.
6 Real data applications
In this section, we apply the proposed methods to two real-data analyses with distinct tasks. The first is a regression setting based on online requests from Stack Exchange and Wikipedia (Danescu-Niculescu-Mizil et al., 2013), where we conduct estimation and inference on the regression coefficient linking perceived politeness to specific linguistic features. The second is an image classification task based on the ImageNet dataset (Russakovsky et al., 2015).
6.1 Politeness of online requests
We study how certain linguistic features affect perceived politeness using the dataset from Danescu-Niculescu-Mizil et al. (2013). The data consist of over ten thousand two-sentence requests collected from Wikipedia and Stack Exchange. Each request was independently annotated by five U.S.-based Amazon Mechanical Turk workers on a 1-25 scale, and the final politeness score is the average of these five ratings. The dataset also includes automatically extracted politeness markers such as gratitude, deference, greetings, sentiment cues, modality choices, and pronominal patterns, obtained by combining dependency parsing with curated lexical resources. In this experiment, we focus on the relationship between politeness and indicative modal features. We select 1000 requests from the dataset and treat the regression coefficient from regressing the politeness score on the indicative modal features as the parameter of interest. Among these requests, we randomly designate 300 as unlabeled data and vary the number of labeled samples $n$ from 50 to 700.
To evaluate the proposed methods, we also obtain predictions for each request from three large language models, GPT-4o, Llama-3-8B, and DeepSeek-V3, by prompting each model to rate the politeness of the given text on a scale of 1-25. For each request, we supply the model with the text under review, relevant background information, and a set of ten independent prompts designed to reduce randomness in the generated scores. The final prediction is computed as the average of the ten outputs. Appendix S9 provides a step-by-step description of the procedure used to generate these predictions.
Figure 3 reports the variation of standard deviations of different methods as the labeled data size increases. The three panels differ only in the implementation of the PPI and PPI++ methods, with each panel using predictions from a different single LLM. While the standard deviations of all methods generally decrease with more labeled data, the PPI and PPI++ estimators do not consistently improve upon the naive approach in terms of efficiency. In contrast, the SADA estimator, which integrates all predictions, consistently outperforms all other methods. Its performance closely tracks that of the best-performing PPI++ estimator based on GPT-4o, highlighting its adaptive behavior. We also plot the oracle SADA estimator, SADAo, which includes the ground truth as one of its predictions. As shown, SADAo performs almost identically to the oracle benchmark obtained by regressing the ground-truth outcome on the covariates using the full data, further confirming the desired adaptivity of the proposed method.